OpenStudy (anonymous):

ln x = 1/3 (ln6 + 2ln2)

OpenStudy (anonymous):

You might notice that: \(\ln x + \ln y = \ln xy\) and \(\ln x^a = a \ln x\) Do you need further help?

OpenStudy (anonymous):

so ln x = 1/3 ln (6*2^2)??

OpenStudy (anonymous):

Yup! :)

OpenStudy (anonymous):

i'm a bit stuck after that. how do i proceed to solve for x?...thank you for responding btw it's very appreciated

OpenStudy (anonymous):

After getting \(ln x = \frac{1}{3} \ln 24\), notice that \(\frac{1}{3} \ln 24=\ln 24^{\frac{1}{3} }\). So, \(x=24^{\frac{1}{3}} \)

OpenStudy (anonymous):

\[\ln x= 1/ \ln 24^3 then solve x\]

OpenStudy (anonymous):

You're welcome :) btw

OpenStudy (anonymous):

I got that at first and I apparently was going about it the right way, my book just gives me a different answer than what i have. I get 2 (sqrt 2) but my book has 4 as the answer. am I doing anything wrong maybe?

OpenStudy (anonymous):

Well, this should shed some light. http://www.wolframalpha.com/input/?i=ln%20x%20%3D%201%2F3%20(ln6%20%2B%202ln2)&t=crmtb01

OpenStudy (anonymous):

that's what I meant I had as my answer, and thanks that did help out :)

OpenStudy (anonymous):

At least you know the book's wrong ;) lol

OpenStudy (anonymous):

yes book ans wrong

OpenStudy (anonymous):

yup lol. thank you guys. good night

OpenStudy (anonymous):

same to u sister.